OpenStudy (anonymous):

let sinA = 3/5 in QI find cosA and tanA

OpenStudy (john_es):

\[\cos A=\sqrt{1-\sin^2 A}=\sqrt{1-(3/5)^2}\] \[\tan A=\frac{\sin A}{\cos A}\]

OpenStudy (usukidoll):

Q1 = quadrant 1? as in first quadrant

OpenStudy (usukidoll):

I would just draw the triangle and use the Pythagorean theorem

OpenStudy (usukidoll):

that would be 16 and the square root is 4

OpenStudy (usukidoll):